Some banks now have biweekly mortgages (that is, with payments every other week). Compare a 20-year, $80,000 loan at 7.5% by finding the payment size and the total interest paid over the life of the loan under each of the following conditions. (Round your answers to the nearest cent.)

1) Payments are monthly, and the rate is 7.5%, compounded monthly. a) payment size $____ b) total interest $_____

2)Payments are biweekly, and the rate is 7.5%, compounded biweekly. (Assume a standard 52-week year.) a) payment size $______ b) total interest $ ______
Answer #1

Loan amount = 80,000; Number of years = 20; annual rate = 7.5%

1). PV = 80,000; N = 20*12 = 240; rate = 7.5%/12 = 0.625%, solve for PMT.

PMT = 644.47 (payment size)

Total amount paid over 20 years = PMT*240 = 644.47*240 = 154,673.89

Total interest paid = total amount paid - loan amount = 154,673.89 - 80,000 = 74,673.89

1). PV = 80,000; N = 20*26 = 520; rate = 7.5%/26 = 0.288%, solve for PMT.

PMT = 297.23 (payment size)

Total amount paid over 20 years = PMT*240 = 297.23*520 = 154,561.99

Total interest paid = total amount paid - loan amount = 154,561.99 - 80,000 = 74,561.99